OpenStudy (anonymous):

suppose the f(x)=3x(5-4x)^3 find an equation for the tangent line to the graph of f at x=1. I dislike these specific types of problems, so if anyone could elaborately explain how to solve these, it would be greatly appreciated.

OpenStudy (amistre64):

f' = r'l + rl' r = 3x; r'=3 l=(5-4x)^3; l' = 3(5-4x)^2 (-4)

OpenStudy (amistre64):

tangent line = f'(1)x -f'(1)1 +f(1)

OpenStudy (amistre64):

f'(1) = 3(1)^3 + 3(1)(3)(-4)(1)^2 = 3 -36 = -33

OpenStudy (amistre64):

tangent = -33x +33+3 y = -33x + 36 maybe?

OpenStudy (amistre64):

do you follow it?

OpenStudy (anonymous):

yeah, it makes sense. I'm always missing one part or another, so I get half or most of the answer correct, just annoying little crap here and there that messes me up.